OpenStudy (anonymous):

Why does √6 x √3 equal 3√2?

geerky42 (geerky42):

\[\sqrt{6}\sqrt{3} \\= \sqrt{3\cdot2}\sqrt{3} \\= \sqrt{3\cdot 2 \cdot 3} \\ = \sqrt{3^2 \cdot 2} \\= \sqrt{3^2}\sqrt{2} \\= \boxed{3\sqrt{2}}\]